OpenStudy (anonymous):

find a function F(x) such that F'(x)=1/x and F(2)=0

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110226092505AAyck8T hope this help

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

dy/dx = 1/x y =ln(x) +C let C = ln(k) y = ln(kx) ln(2k) = 0 2k = 1 k = 1/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

is my answer clear? tell me if what i have written is confusing

OpenStudy (anonymous):

no not at all

OpenStudy (anonymous):

i understood it clearly

OpenStudy (anonymous):

thanks a lot